OpenStudy (specek18):

REALLY NEED HELP WILL GIVE MEDALS FOR CORRECT ANSWER...PLEASE!!! factor the polynomial using the pattern X^2+(a+b)x+ab=(x+a)(x+b) that's the pattern here's the polynomial x^2+3x-18

OpenStudy (specek18):

@Abhisar do you know/ can you help? @Brostep0s @Coolsector @Directrix @Donblue22 @CausticSyndicalist @Escobar17 @frizzank @Fazeelayaz @danish071996 @quin100 @TuringTest @KamiBug @kmorgan

OpenStudy (anonymous):

so we have a+b = 3 a*b=-18 can you find a and b ?

OpenStudy (specek18):

working on it. one moment:) @Coolsector

OpenStudy (anonymous):

do you understand how i got those 2 equations ?

OpenStudy (specek18):

okay i thought i did at first because i know i have to have a (a+b) and an (a*b) that equal the same thing, but i am having trouble finding it...

OpenStudy (anonymous):

the general form is : X^2+(a+b)x+ab and our poly is : x^2+3x-18 so i look at the terms you see in our poly 3 is multiplying x and in the form it is (a+b) so : a+b =3 the free term in the form is ab and in our poly is -18 so ab = -18

OpenStudy (specek18):

okay so is -18 the final answer? i still don't get what adds & multiplies to get -18...thank you so much for your patience with me!!!

OpenStudy (anonymous):

no, we have two equations : a+b=3 and ab=-18 i explained how i got them.. are you having trouble solving them ?

OpenStudy (specek18):

yeah...i'm sorry i have such a hard time pin pointing where i'm struggling but i think that would be a great help if you could help me solve them?

OpenStudy (anonymous):

sure, but I want you to understand how I got the equations (since it is the first step towards the solution). so our equations : a+b=3 ab = -18 we can use the first one and get: b = 3-a now plug it into the second one and get: a(3-a) = -18 a^2-3a-18=0 solving this quad equation give us : a=6 , -3 plug it into our b=3-a : b = -3,6 so you can choose a=6 , b=-3 or the opposite . in any case you get: (x-6)(x+3)

OpenStudy (specek18):

so it just all disappeared....on my school:(((( i don't know why but if i get it to reload i'll contact you. CANNOT THANK YOU ENOUGH!!!

OpenStudy (anonymous):

ok sure whatever you need